Construction. A garden area is 30 ft long and 20 ft wide. A path of uniform width is set around the edge. If the remaining garden area is 400 ft^2, what is the width of the path?
Thanks:)

Answer by checkley71(8403)

THE DIMENTIONS OF THE REMAINING GARDEN IS (30-2X)*(20-2X)=400 OR
600-40X-60X+4X~2=400 OR 600-400-100X+4X~2 OR 4X~2-100X+200 OR X~2-25X+50
USING THE QUADRATIC EQUATION WE GET X=2.19 & X=22.8 THUS THE DIMENTIONS OF THE INNER GARDEN IS 30-4.38=25.62 & 20-4.38=15.62 OR 25.62*15.62=400
THUS THE PATH IS ~ 2.19 FEET WIDE